OpenStudy (anonymous):

okay so the derivative of f(x)=cscx is f'(x)=-csc x cot x. Now can someone help me find the x-values where the first derivative is zero or undefined

OpenStudy (zzr0ck3r):

you have -csc(x)cot(x) = -1/sin(x) * cos(x)/sin(x) now -1/sin(x) is never equal to 0, so the only way the whole thing will be zero is if cos(x)/sin(x) = 0 this will only happen when cos(x) = 0 this happens when x = pi/2 + pi* k

OpenStudy (zzr0ck3r):

that should say \(x=\frac{\pi}{2}\pm\pi*k\) where \(k\in\mathbb{Z}\)

OpenStudy (anonymous):

thanks zzr0ck3r but what about undefined?

OpenStudy (anonymous):

y' = -cos x/sin^2 x y' = 0 when cos x = 0-> x = Pi/2 and 3Pi/2 (0 < Pi <2Pi) Y' undefined when sin x = 0-> x = pi , or 2Pi

OpenStudy (zzr0ck3r):

its undefined when -1/sin(x) is undefined and when cot(x) = cos(x)/sin(x) is undefined this happens when sin(x) = 0 and that is when \(x=2\pi*k\) where \(k\in \mathbb{Z}\)

OpenStudy (zzr0ck3r):

sorry * and that is when \(x=\pi*k\) where \(k\in \mathbb{Z}\)

OpenStudy (zzr0ck3r):

@thu1935 its undefined at infinite points

OpenStudy (jdoe0001):

\(\bf -\cfrac{1}{sin(x)=0}\cdot \cfrac{cos(x)}{sin(x)=0}\implies -\cfrac{1}{0}\cdot \cfrac{cos(x)}{0}\)

OpenStudy (zzr0ck3r):

holy abuse in notation batman

OpenStudy (jdoe0001):

heheh

OpenStudy (zzr0ck3r):

:P

OpenStudy (jdoe0001):

http://www.wolframalpha.com/input/?i=batman+curve

OpenStudy (zzr0ck3r):

wow crazy